OpenStudy (anonymous):

Define product rule and give me three examples.. its for a test i have on Monday my teacher said that i can write the definition and three examples

OpenStudy (hba):

The probability of two independent events occurring simultaneously is the product of the individual probabilities. The rule stating that the probability of the occurrence of independent events is the product of their separate probabilities.

OpenStudy (anonymous):

Okay what are three examples?(:

OpenStudy (anonymous):

Oh and its the rule in exponets.

hero (hero):

Yeah, there's like three different product rules. One for exponents, one for calculus, and one for probability.

hero (hero):

But for exponents \(a \times a = a^2\) for example

OpenStudy (lgbasallote):

i dont think it's asking about "three types of product rule" because there are more than three....

OpenStudy (lgbasallote):

the product rule of exponents though is \[\Large a^m \times a^n = a^{m+n}\] an example would be \[\Large x^3 \times x^4 = x^{3+4} = x^7\]